OpenStudy (anonymous):

URGENT:FUNDAMENTAL THEOREM OF SIMILARITY!!! A regular hexagon with a side length of 4 cm has an exact area of 24 square root 3 cm squared. What is the exact area of similar hexagon with a side length of 14 cm?

OpenStudy (mertsj):

Just remember this: The ratio of the areas is the square of the ratio of the sides.

Directrix (directrix):

If two polygons are similar, the square of the scale factor of the two polygons is equal to the ratio of any two corresponding area measurements of the solids. (4/14)² = (24√3) / A where A is the area of the polygon with side 14. (2/7)² = (24√3) / A 4/49 = (24√3) / A 4A = 49 * (24√3) Sove for A --> @haley321